1. Name one point(other than the y-intercept) that is on the line with the given equation y= -2x + 8?
2. What is the y-intercept of the line with an equation of -2x + 5y - 10=0?
3. write the slope-intercept form of 8x + 4y=32?
4. Find the slope given the ordered pairs (3,-2), (12,10)?
5. The slope of the linear equation 4x + 2y=10?
6. sketch the graph of the equation y= 3/5x - 2?

Answer by lynnlo(4176)

(1)slope
(2)(0,2)
(3)y=-2x+8
(4)m=4/3
(5)m=-2
(6)you should be able to graph y=3/5x-2